# Interpretable Concept Bottlenecks to Align Reinforcement Learning Agents

### v2

@@ 3 Experimental Evaluations
To validate these findings further, we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (referred to as NoEnemy cf. Fig. 1). We reevaluate the deep and SCoBot agents that were trained on the original Pong environment on NoEnemy and observe catastrophic performance drops in HN scores (cf. Fig. 3) at a level of random policies for both types of agents. Evaluations with different DQN agents lead to similar performance drops (cf. App. A.7). These drops are particularly striking as initial importance maps produced by PPO and DQN agents highlight all $3$ moving objects (i.e. the player, ball, and enemy) as relevant for their decisions (cf. Fig.1 top left and App. A.7) which is aligned with findings on importance maps of (Weitkamp et al., 2018). These maps suggest that the deep agents had based its decisions on right reasons. Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and granularity of such post-hoc explainability methods in the context of RL. Specifically, they highlight the importance of inherently explainable RL models on the level of concepts, as provided via SCoBots, to identify such potential issues. In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottleneck structure.
@@
Realigning SCoBots: For mitigating goal misalignment issues of the SCoBot agents trained on Pong, we can simply remove (prune) the enemy from the set of considered objects ($\overline{\Omega}$). By doing so, the enemy cannot be used for the action selection process. This leads to guided SCoBots that are able to play Pong and its NoEnemy version (cf. guided SCoBot in Fig 3 (center)). Furthermore, this shows that playing Pong without observing the enemy is achievable, by simply returning vertical shots, difficult for the enemy to catch.
@@
$\begin{split}R^{\text{exp}}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}&-D(Player,o)^{t-1}\\[-8.61108pt] &+V(Player).y.\end{split}$ (5)
These rewards are further detailed in App. A.8. With this, we obtain guided SCoBot agents that reach human scores, whereas the deep and unguided SCoBot agents perform worse than random (cf. Fig. 3 (left)).
Note that providing additional reward signals, as done above, obviously also allows to mitigate reward sparsity, as we illustrate on Pong (cf. App. A.8.3 for a detailled explanation).
@@ 4 Limitations
In our evaluations, we made use of the integrated object extractor ($\omega$) of OCAtari which provides us with ground truth data. Such extractors can however also be optimized using supervised (Redmon et al., 2016) or self-supervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods. Furthermore, our set of relational functions, $\mathcal{F}$, is predefined and fixed. To eliminate the reliance on expert users, this set may alternatively be obtained using the supervision from large language models (Wu et al., 2023), via prior task-knowledge integration (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b). Other environments require additional information extraction processes. E.g. in MsPacman an agent must navigate a maze. Providing concept representations for such environments is an important step for future work. Finally, during our revision assessments, we presuppose that the user has adequate knowledge and positive intentions.
@@ 5 Related Work
The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as Lampert et al. (2009) and Kumar et al. (2009). A first systematic study of CBMs and their variations was delivered by Koh et al. (2020) and Stammer et al. (2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output. Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Stammer et al., 2022; Steinmann et al., 2023; Sawada & Nakamura, 2022; Marconato et al., 2022; Lage & Doshi-Velez, 2020), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$. Since the initial works on CBMs they have found utilization in several applications. Antognini & Faltings (2021) e.g. apply CBMs to text sentiment classification. However, these works consider single object concept representations and focus on supervised learning.
@@
In fact, CBMs have found their way into RL only to a limited extent. Zabounidis et al. (2023) and Grupen et al. (2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance. Zabounidis et al. (2023) further report better training stability and reduced sample complexity. Their extension includes an optional “residual” layer which passes additional, latent information to the action selector part. SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations. Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, Guan et al. (2023) allow the user to design a reward function based on higher-level properties that occur over a period of time. Lastly, SCoBots not only separate state representation learning from policy search, as done by Cuccu et al. (2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
@@
Explainable RL (XRL) is an extensively surveyed XAI subfield (Milani et al., 2023; Dazeley et al., 2022; Krajna et al., 2022) with a wide range of yet unsolved issues (Vouros, 2022). Milani et al. (2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior. Based on this, SCoBots extract relations from low-level features making high-level information available to explanations and thereby support feature importance methods. According to (Qing et al., 2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining).
@@
The misalignment problem is an RL instantiation of the shortcut learning problem, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023), prototype-based models (Bontempelli et al., 2023) and RL approaches (di Langosco et al., 2022). A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold & Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022). Nahian et al. (2021) ethically align agents by introducing a second reward signal. In comparison, SCoBots aid to resolve RL specific issues such as the misalignment problem through inherent interpretability.
@@ 6 Conclusion
Avenues for future research are incorporating a high level action bottleneck (Bacon et al., 2017). One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by (Wu et al., 2023). Additionally, we are considering the use of shallower decision trees (Broelemann & Kasneci, 2019). An interesting research question is how far the task reward signal can aid in learning games object-centric representations (Delfosse et al., 2023c) in the first place.
@@ A.1 Atari games are most common set of environments
We here show that the Atari games from the Arcade Learning Environment (Bellemare et al., 2012) is the most used benchmark to test reinforcement learning agents.
Figure 5: The Atari Learning Environments is more used in scientific research than the next 88 other benchmarks together. Graph borrowed from (Delfosse et al., 2023a).
@@ A.2 Detailed numerical results
All agents are trained on gymnasium’s atari learning environments using $8$ actors and $3$ training seeds ([0,16,32]+[0,16,32]+rank). Each training seed’s performance is evaluated every 500500k frames on $4$ differently seeded (42+42+training seed) environments for $8$ episodes each. After training, the best performing checkpoint is then ultimately evaluated on $4$ seeded ($123,456,789,1011$) test environments. The final return is determined by averaging the return over $5$ episodes per training-test seed pair and every training seed for the respective environment. We use deterministic actions for both evaluation and testing stages.
@@
SCoBots
v5
SCoBots
guided-v5
PPO
ours-v5
PPO
ours-v4
PPO
Schulman et al.-v4
@@
Asterix $5080\mbox{\scriptsize$\pm 614.8$}$ $5043.3\mbox{\scriptsize$\pm 729.9$}$ $2126.7\mbox{\scriptsize$\pm 148.6$}$ $10433.3\mbox{\scriptsize$\pm 2772.9$}$ $4532$ $210$ $8503$
@@
Table 1: ScoBots and guided ScoBots obtain similar results than deep agents averaged over 33 seeded runs. Neural refers to the agent that use a CNN instead of our ICB layers. We added the results reported in the original paper (Schulman et al., 2017) (original), which are on par with ours, as well as ones from a Random baseline and Humans (from (van Hasselt et al., 2016)). Our agents have been trained with only 2020M frames.
@@ A.3 Learning curves
Figure 6: Training Return over frames ($\times 10^{6}$) seen for   guided SCoBots,   SCoBots and   neural agents.
@@ A.5 Hyperparameters and experimental details
All Experiments were run on a AMD Ryzen 7 processor, 64GB of RAM and one NVIDIA GeForce RTX 2080 Ti GPU. Training a (i)SCoBots on $20\text{M}$ frames and using $8$ actors takes approximately $4$ hours. We use the same PPO hyperparameters as the Schulman et al. (2017) agents that learned to master the games. For the Adam optimizer, SCoBots start with a slightly increased learning rate of $1\times 10^{-3}$ (compared to $2.5\times 10^{-4}$). The PPO implementation used and the respective MLP hyperparameters are based on stable-baselines3 Raffin et al. (2021). SCoBots have the same PPO hyperparameter values as deep agents but use MLPs ($2\times 64$) with ReLU activation functions as policy and value networks. Deep agents use CnnPolicy in stable-baselines3 as their policy value network architecture, which aligns with the initial baseline implementation of PPO. Further, we normalize the advantages in our experiments, since it showed only beneficial effects on learning. This is the default setting in the stable-baseline3 implementation.
The Atari environment version used in gymnasium is NoFrameskip-v4 for agents reproducing the reported PPO results, and v5 for SCoBots and neural. v5 defines a deterministic skipping of 5 frames per action taken and sets the probability to repeat the last action taken to 0.25. This is aligned with recommended best practices by Machado et al. (2018). The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation444github.com/openai/baselines. This includes frame skipping of 4 and reward clipping. Frame stacking is not used. SCoBots are not trained on a clipped reward signal. For comparability, the neural agents we compare SCoBots to, are not as well. A list of all hyperparameter values used is provided in Table 2.
@@ A.5.1 The properties and features used for SCoBots
Properties class NAME object class (e.g. ”Agent”, ”Ball”, ”Ghost”)
@@ A.6 SCoBots policies: formal definitions
The set $\mathcal{S}$ denotes the problem-specific set of raw environment states as defined by the RL problem, e.g., the space of RGB images $0,255]^{512\times 512\times 3}$. Similarly, $\mathcal{A}$ represents the action space, e.g., “move right” or “jump.”

### v4

@@ 3 Experimental Evaluations
To validate these findings further, we perform evaluations on $2$ modified Pong environments in which (i) the enemy is invisible, yet playing (NoEnemy, cf. Fig. 1), and on one environment where the enemy is not moving after returning the ball (LazyEnemy). We reevaluate the deep and SCoBot agents that were trained on the original Pong environment on NoEnemy and observe catastrophic performance drops in HN scores (cf. Fig. 3) at a level of random policies for both types of agents. Evaluations with different DQN agents lead to similar performance drops (cf. App. A.7). These drops are particularly striking as initial importance maps produced by PPO and DQN agents highlight all $3$ moving objects (i.e. the player, ball, and enemy) as relevant for their decisions (cf. Fig.1 top left and App. A.7), aligned with findings on importance maps of Weitkamp et al. (2018). These maps suggest that the deep agents had based its decisions on right reasons. Our novel results on the NoEnemy and LazyEnemy environments, however, greatly calls to question the faithfulness and granularity of such post-hoc explainability methods. They highlight the importance of inherently transparent RL models on the level of concepts, as provided via SCoBots, to identify such potential issues. In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottlenecks.
@@
Realigning SCoBots: To mitigate the goal misalignment issues of the SCoBots trained on Pong, we simply remove (prune) the enemy from the set of considered objects ($\overline{\Omega}$). Thus, the enemy cannot be used for the action selection process. This leads to SCoBots that are able to play Pong and its NoEnemy version (cf. SCoBot in Fig 3 (center)). Furthermore, this shows that playing Pong without observing the enemy is achievable, by simply returning vertical shots, difficult for the enemy to catch.
@@
$R^{\text{exp}}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}-D(Player,o)^{t-1}+V(Player).y.$ (2)
These rewards are further detailed in App. A.8. They allow guided SCoBots to reach human scores, whereas the deep and unguided SCoBots perform worse than random (cf. Fig. 3 (left)). Note that providing additional reward signals, as done above, obviously also allows to mitigate reward sparsity, as we illustrate on Pong (cf. App. A.8.3 for a detailed explanation).
Figure 5: SCoBots can learn with noisy object detectors, transparent SCoBots rely on relations. Final human normalized scores (with stds) comparing SCoBots and the object-centric neural baselines (NN-SCoBots), with and without relations. We also provide the scores of NN-SCoBots that learned on noisy environments. The noise only noticeably affects the agents on Kangaroo. Ablating the relations is harmless on NN-SCobots, as neural networks can recompute them, but impacts SCoBots performances on 66 games. (For better visualization, we used a human score of $100$ in Boxing.)
Object-centric agents can learn with imperfect object extractors (Q4).
To test object-centric agents’ ability to work in more realistic environments, we have tested if they can learn viable policies with imperfect object extraction methods. Based on Delfosse et al. (2023b)’s detection results, we added a $5\%$ misdetection probability and a Gaussian noise (of 0 mean and $3$ pixels of standard deviation on each axis). As depicted in Figure 5, NN-SCoBots (i.e. neural network based object-centric agents, using relations) learn comparable policies on all games but Kangaroo, demonstrating their ability to mitigate suboptimal object extraction methods. Implementing robustness techniques, such as Kalman filters, would help to further push the performances of such agents.
Transparent SCoBots benefit from explicit relations (Q5).
While the ablation of the relation extractor only has a significant impact on the neural-based object-centric agents on $1$ out of $9$ games (Seaquest), it deteriorates the decision tree based SCoBots on $6$ environments. This is due to the fact that relations can be implicitly recomputed within a neural network, but not within decision trees. As shown by Kohler et al. (2024), the use of the distance relation (on a specific axis) allows for performing agents with compact decision tree-based policies. Furthermore, even if some relations can be rediscovered and implicitly encoded within the decision trees, the lack of explicit relational representations can reduce the interpretability of the agents, and will prevent the experts from using these within their guidance.
@@ 4 Limitations
On the use of OCAtari. To limit our resource consumption, we made use of the quasi-perfect object extractor ($\omega$) of OCAtari, which efficiently extract objects from the RAM. We added an ablation to simulate potential imperfect detection capabilities of Atari object extraction methods. Such extractors can be optimized using supervised (Redmon et al., 2016; Locatello et al., 2020) or self-supervised (Lin et al., 2020; Delfosse et al., 2023b) object detection methods. This last work showcase that unsupervised object extraction methods can replace OCAtari at test time, however leading to performance drops. Grandien et al. (2024) have further showcase training RL agents with such pretrained object extractors further improve the object-centric RL agents performances.
The limit of object-centricity. Other environments require additional information extraction processes. E.g. in MsPacman, an agent must navigate a maze. Extending the concept representations to cover such concepts in maze or platform environments is an important step for future work. Other representations could allow for the integration of e.g. path finding methods such as the A* algorithm.
Training time of SCoBots. Thanks to the efficient OCAtari object extraction, SCoBots required in average $7.5$ hours of training time, while deep agents needed $10.8$ hours (cf. Appendix A.9). All SCoBots variations require less training time compared to deep agents on environments with few objects (e.g. Boxing, Pong, Tennis). Currently, at every step, the concept bottleneck values are calculated sequentially in a single process on the CPU, leaving significant room for training time improvement by optimizing these computations (bringing them to GPUs). Thus, in environments with many objects, e.g. Kangaroo, the training time of unguided SCoBots exceeded that of its deep agent counterpart. Attention on relevant objects could be used to further save computational resources.
@@ 5 Related Work
The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as Lampert et al. (2009) and Kumar et al. (2009). A first systematic study of CBMs was delivered by Koh et al. (2020), followed by Stammer et al. (2021), who described CBMs as two-stage models that first computes intermediate representations used for the final task output. Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Lage and Doshi-Velez, 2020; Stammer et al., 2022; Sawada and Nakamura, 2022; Marconato et al., 2022; Steinmann et al., 2023; Stammer et al., 2024), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$. Since the initial works on CBMs they have found utilization in several applications. Antognini and Faltings (2021) e.g. apply CBMs to text sentiment classification, and Kraus et al. (2024) to time series’ analysis. However, these works consider single object concept representations and focus on supervised learning. The ability of users to revise concepts and decisions has also been parallelly shown by Friedrich et al. (2023).
@@
In fact, CBMs have found their way into RL only to a limited extent. Zabounidis et al. (2023) and Grupen et al. (2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance. Zabounidis et al. (2023) further report better training stability and reduced sample complexity. Their extension includes an optional “residual" layer which passes additional, latent information to the action selector part. SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations. Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, Guan et al. (2023) allow the user to design a reward function based on higher-level properties that occur over a period of time. Lastly, SCoBots not only separate state representation learning from policy search, as done by Cuccu et al. (2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
@@
Explainable RL (XRL) is an extensively surveyed XAI field (Dazeley et al., 2022; Krajna et al., 2022; Milani et al., 2023) with a wide range of unsolved issues (Vouros, 2022). Milani et al. (2023) introduce a taxonomy for XRL methods, with: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present which past experience or MDP components affect the policy, and (3) policy-level methods, describing the agent’s long-term behavior. Based on this, SCoBots extract relations from low-level features, making high-level information available to explanations and thereby support feature importance methods. According to Qing et al. (2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining). Other XRL methods rely on LLM to explain the policy (Luo et al., 2024) decision trees (Fuhrer et al., 2024; Marton et al., 2024), logic (Jiang and Luo, 2019; Kimura et al., 2021; Delfosse et al., 2023a; Sha et al., 2024) or programs (Verma et al., 2018; Trivedi et al., 2021; Cao et al., 2022; Kohler et al., 2024; Wüst et al., 2024) to encode transparent policies.To overcome the potentially unavailable concepts necessary to learn symbolic policies, Shindo et al. (2024) learn a mixture of symbolic and neural policies. Finally, concepts have further been used to derive reward from context using LLMs, as we did for Kangaroo and Skiing (Kwon et al., 2023; Kaufmann et al., 2024; Wu, 2024; Shen et al., 2024).
@@
The misalignment problem is an RL instantiation of the shortcut learning problem, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023), prototype-based models (Bontempelli et al., 2023) and RL approaches (di Langosco et al., 2022). A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold and Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022). Nahian et al. (2021) ethically align agents by introducing a second reward signal. In comparison, SCoBots aid to resolve RL specific issues such as the misalignment problem through inherent interpretability.
@@ 6 Conclusion
Avenues for future research are incorporating a high level action bottleneck (Bacon et al., 2017). One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by Wu et al. (2024). Additionally, we are considering the use of shallower decision trees (Broelemann and Kasneci, 2019). An interesting research question is how far the task reward signal can aid in learning games object-centric representations (Delfosse et al., 2023b) in the first place.
@@
Acknowledgments
This work has benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, their joint support of the National Research Center for Applied Cybersecurity ATHENE, via the “SenPai: XReLeaS” project. It has also received support of Hessian.AI, as well as the Hessian research priority program LOEWE within the project WhiteBox, and the EU-funded “TANGO” project (EU Horizon 2023, GA No 57100431). The authors also would like to thank Stefan Lichtenstein, Raban Emunds for their help on refactoring the code, and Dwarak Vittal for his initial contributions.
@@ A.1 Atari games are most common set of environments
Figure 6: The Atari Learning Environments is more used in scientific research than the next 88 other benchmarks together. Graph borrowed from [Delfosse et al., 2024b].
We here show that the Atari games from the Arcade Learning Environment [Bellemare et al., 2012] is the most used benchmark to test reinforcement learning agents.
@@ A.2 Detailed numerical results
All agents are trained on gymnasium’s ALE using $8$ actors and $3$ training seeds ([0,16,32]+[0,16,32]+rank). Each training seed’s performance is evaluated every 500500k frames on $4$ differently seeded (42+42+training seed) environments for $8$ episodes each. After training, the best performing checkpoint is then ultimately evaluated on $4$ seeded ($123,456,789,1011$) test environments. The final return is determined by averaging the return over $5$ episodes per training-test seed and every training seed for the respective environment. We use deterministic actions for both evaluation and testing stages.
@@
SCoBots-v5
NoGuidance
SCoBots-v5
PPO-v5
ours
PPO-v4
ours
PPO-v4
Schulman et al.
@@
Asterix $5080\mbox{\scriptsize$\pm 614.8$}$ $5043.3\mbox{\scriptsize$\pm 729.9$}$ $2126.7\mbox{\scriptsize$\pm 148.6$}$ $10433\mbox{\scriptsize$\pm 2773$}$ $4532$ $210$ $8503$
@@
Table 1: ScoBots and guided ScoBots obtain similar results than deep agents averaged over 33 seeded runs. Neural refers to the agent that use a CNN instead of our ICB layers. We added the results reported in the original paper Schulman et al. [2017] (original), which are on par with ours, as well as ones from a Random baseline and Humans (from van Hasselt et al. [2016]). Our agents have been trained with only 2020M frames.
@@ A.3 Learning curves
Figure 7: Training Return over frames ($\times 10^{6}$) seen for   SCoBots,   SCoBots w/o guidance and   neural agents.
@@ A.5 Hyperparameters and experimental details
All Experiments were run on a AMD Ryzen 7 processor, 64GB of RAM and one NVIDIA GeForce RTX 2080 Ti GPU. Training a SCoBot on $20\text{M}$ frames with $8$ actors takes approximately $8$ hours. We use the same PPO hyperparameters as the Schulman et al. [2017] agents that learned to master the games. For the Adam optimizer, SCoBots start with a slightly increased learning rate of $1\times 10^{-3}$ (compared to $2.5\times 10^{-4}$). The PPO implementation used and the respective MLP hyperparameters are based on stable-baselines3 Raffin et al. [2021]. SCoBots have the same PPO hyperparameter values as deep agents but use MLPs ($2\times 64$) with ReLU activation functions as policy and value networks. Deep agents use CnnPolicy in stable-baselines3 as their policy value network architecture, which aligns with the initial baseline implementation of PPO. Further, we normalize the advantages in our experiments, since it showed only beneficial effects on learning. This is the default setting in the stable-baseline3 implementation.
The Atari environment version used in gymnasium is NoFrameskip-v4 for agents reproducing the reported PPO results, and v5 for SCoBots and neural. v5 defines a deterministic skipping of 5 frames per action taken and sets the probability to repeat the last action taken to 0.25. This is aligned with recommended best practices by Machado et al. [2018]. The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation444github.com/openai/baselines. This includes frame skipping of 4 and reward clipping. Frame stacking is not used. SCoBots are not trained on a clipped reward signal. For comparability, the neural agents we compare SCoBots to, are not as well. A list of all hyperparameter values used is provided in Table 2.
@@ A.5.1 The properties and features used for SCoBots
Properties class NAME object class (e.g. "Agent", "Ball", "Ghost")
@@ A.6 SCoBots policies: formal definitions
The set $\mathcal{S}$ denotes the problem-specific set of raw environment states as defined by the RL problem, e.g., the space of RGB images $0,255]^{512\times 512\times 3}$. Similarly, $\mathcal{A}$ represents the action space, e.g., “move right" or “jump."